OpenStudy (anonymous):

Suppose $10,000 is divided into two bank accounts. One account pays 3% simple interest per year and the other pays 2%. After one year there is a total of $240 in interest between the two accounts. How much was invested into the bank account that pays 3% simple interest (rounded to the nearest cent)? $5,000 $4,000 $6,000 $10,000 @gorv

OpenStudy (gorv):

let first one investment =x

OpenStudy (gorv):

second investment =10000-x

OpenStudy (anonymous):

9700

OpenStudy (gorv):

x*3/100+(10000-x)*2/100=240

OpenStudy (gorv):

multiply both side by 100\ 3x+(10000-x)*2=24000 3x+20000-2x=24000 x+20000=24000 subtract 20000 from both side

OpenStudy (gorv):

x=4000 at rate of 3%

OpenStudy (anonymous):

thanks

OpenStudy (gorv):

@DUDE..LOOK..ITS..SGTC

OpenStudy (gorv):

welcome:)